A quarterly gas bill is given by the formulae c = S + nx
where C is the total cost of the bill in pence
S is the standing charge in pence
n is the number of units used
x pence is the cost per unit.
(a)calculate the value of C when S = 860, n = 152 and x =7
(b)Rewrite the formula so that it is written in the form n =...

Answer by stanbon(75887)

Put the numbers in place of the letters to get:
C = 860+(152)(7)
C= 1924
-----------------
C= S +nx
Substract "S" from both sides to get:
nx = C-S
divide by "x" to get:
n = (C-S)/x
Cheers,
Stan H.